A storage battery of emf 12V and internal resistance 0.5 ohm, is to be charged by a 120V D.C supply of negligible internal resistance. What resistance is required in the circuit for the charging current to be 3A? 
Calculate:
i)How much chemical energy from the dc source is converted to electrical energy?
ii)How much electrical energy is converted to chemical energy in the storage battery?

Grade:12

2 Answers

Arun
4 years ago

 

By applying Kirchoff's voltage law , we get,   3R + 12 + (3×0.5) = 120 
 
From abov eqn., we get R = 35.5 Ω
 
If t is charging time in seconds, electrical energy converted into chemical energy in storage battery = ( 12 × 3 × t ) J = 36t J
 
If t is charging time in seconds, chemical energy converted into electrical energy in dc source = ( 120 × 3 × t ) J = 360t J
Vikas TU
4 years ago
Dear student 
The Resistance value is incorrect 
Applying Kirchoff voltage rule we get
net emf =     (120-12)=117 V , net resistance =( R+r)   and net current =3A
​now 117=i xRnet
117/3 =r+R
39-0.5 =R=38.5 Ω
Hope this will help